OpenStudy (anonymous):

Suppose you have a drawer full of white, black, and yellow pairs of socks. If the probability of picking a white pair of socks is 4/9, and the probability of picking a black pair of socks is 7/18, what is the probability of picking a yellow pair of socks

OpenStudy (acxbox22):

4/9 + 7/18 + x =18/18 8/18 +7/18+x=18/18 15/18 + x =18/18 x= 3/18 x=1/6 so the probability of picking a yellow pair is 1/6

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you help me with another

OpenStudy (triciaal):

total is always 1 white + black + yellow = 1 4/9 + 7/18 + x = 1 where x is the probabilty for yellow solve for x